# Experimental Design for Active Transductive Inference in Large Language Models

### v1

@@ B.1 Proof of Lemma 1
Now note that $\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}$ is a positive semi-definite matrix for any $\mathbf{x}_{j}$ and any positive semi-definite matrix $\mathbf{A}$. As a result, $\mathbf{x}_{*}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*}\geq 0$ holds for any $\mathbf{x}_{*}$ and thus
$f(S_{+})=\mathbf{x}_{*}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*}^{\top}-\mathbf{x}_{*}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*}\leq\mathbf{x}_{*}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*}^{\top}=f(S)\,.$
@@
B.2 Proof of Theorem 4
The proof is under the assumption that points can be partitioned as $\mathcal{X}=S\cup\bar{S}$. The set $S$ represents points that are close to $\mathbf{x}_{*}$. It is convex and defined by $\alpha\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\geq\alpha$ for all $\mathbf{x},\mathbf{y}\in S$. The set $\bar{S}$ represents points that are not close to $\mathbf{x}_{*}$. It is defined $\beta\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta$ for all $\mathbf{x}\in S$ and $\mathbf{y}\in\bar{S}$. The inverse of the covariance matrix after $t-1$ observations is
$\widehat{\mathbf{\Sigma}}_{t}^{-1}=\mathbf{\Lambda}_{t}=\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{\top}=\sum_{i=1}^{d}\lambda_{t,i}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\,.$
The latter is the eigendecompositon of $\mathbf{\Lambda}_{t}$, where $\lambda_{t,i}$ is the $i$-th largest eigenvalue and $\mathbf{v}_{t,i}$ is the corresponding eigenvector. Note that $\mathbf{\Lambda}_{t}^{-1}=\sum_{i=1}^{d}\lambda_{t,i}^{-1}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}$. To simplify exposition, we assume that all points have unit length. We analyze the eigenvalues of $\mathbf{\Lambda}_{t}$ first.
@@ B.2 Proof of Theorem 4
For all $i\in[d]$, $1\leq\lambda_{t,i}\leq t$. Moreover, let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\lambda_{t,1}\geq\alpha^{2}(t-1)+1$.
@@
The first claims follows directly from the definition of $\mathbf{\Lambda}_{t}$ and that $\|\mathbf{x}_{\ell}\|_{2}\leq 1$. The second claim is proved using the definition of the maximum eigenvalue,
$\lambda_{t,1}=\mathbf{v}_{t,1}^{\top}\mathbf{\Lambda}_{t}\mathbf{v}_{t,1}\geq\mathbf{x}_{*}^{\top}\left(\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{\top}\right)\mathbf{x}_{*}\geq\alpha^{2}(t-1)+1\,.$
The last inequality follows from $\mathbf{x}_{*}^{\top}\mathbf{x}_{\ell}\geq\alpha$. This completes the proof. ∎
@@
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\mathbf{v}_{t,1}\in S$. Moreover, let $\beta\geq 1-\alpha^{2}$. Then $\mathbf{v}_{t,i}\in\bar{S}$ for all $i\geq 2$.
@@
Since all $\mathbf{x}_{\ell}\in S$ and $S$ is a convex set, $\mathbf{v}_{t,1}\in S$. Now take any $\mathbf{x}\in S$ and $i\geq 2$, and note that
$(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}\leq\sum_{i=2}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1-(\mathbf{x}^{\top}\mathbf{v}_{t,1})^{2}\leq 1-\alpha^{2}\,.$
We use that $\sum_{i=1}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1$ and $\mathbf{x}^{\top}\mathbf{v}_{t,1}\geq\alpha$. Therefore, when $\beta\geq 1-\alpha^{2}$, we have $\mathbf{v}_{t,i}\in\bar{S}$ for all $i\geq 2$. ∎
Our analysis has two parts. First, we bound the approximation error under the assumption that $\mathbf{x}_{\ell}\in S$ holds for all $\ell\in[T]$. Second, we show how to choose $\alpha$ and $\beta$ to guarantee this. We start with the approximation error.
@@
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[T]$. Then
$\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*}\leq\frac{1}{\alpha^{2}T+1}+(1-\alpha^{2})\,.$
@@
$\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*}=\sum_{i=1}^{d}\lambda_{T+1,i}^{-1}\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i}\mathbf{v}_{T+1,i}^{\top}\mathbf{x}_{*}\leq\frac{(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,1})^{2}}{\alpha^{2}T+1}+\sum_{i=2}^{d}(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i})^{2}\,.$
The inequality uses lower bounds in Lemma 6. Then we apply $\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,1}\leq 1$ and $\sum_{i=2}^{d}(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i})^{2}\leq 1-\alpha^{2}$. ∎
Now we prove by induction that $\mathbf{x}_{\ell}\in S$ holds for all $\ell\in[T]$.
@@
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Suppose that $t\leq\frac{\alpha^{2}}{(\beta+\sqrt{2})\beta d}$. Then $\mathbf{x}_{t}\in S$.
@@
Our algorithm chooses a point $x_{t}\in S$ when
$\frac{\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}}{1+\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}}\geq\frac{\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}}{1+\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}}$
holds for any $\mathbf{x}\in S$ and $\mathbf{y}\in\bar{S}$. Since $0\leq\mathbf{v}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{v}\leq 1$ for any $\|\mathbf{v}\|_{2}\leq 1$, the above event occurs when
$(\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x})^{2}=\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}\geq 2\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}=2(\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y})^{2}\,.$
@@
$|\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}|\leq\sum_{i=1}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{y}|\leq\beta d\,.$
Here we use $\lambda_{t,i}\geq 1$ (Lemma 6), and that $\mathbf{v}_{t,1}^{\top}\mathbf{y}\leq\beta$ and $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta$ when $i\geq 2$.
@@
$|\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}|\geq\lambda_{t,1}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\mathbf{v}_{t,1}^{\top}\mathbf{x}|-\sum_{i=2}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{x}|\geq\frac{\alpha^{2}}{t}-\beta^{2}d\,.$
To bound the first term, we use $\lambda_{t,1}\leq t$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\geq\alpha$ and $\mathbf{v}_{t,1}^{\top}\mathbf{x}\geq\alpha$. To bound the second term, we use $\lambda_{t,i}\geq 1$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta$ and $\mathbf{v}_{t,i}^{\top}\mathbf{x}\leq\beta$.
Now we chain all inequalities and get that our algorithm chooses a point $x_{t}\in S$ when
$t\leq\frac{\alpha^{2}}{(\beta+\sqrt{2})\beta d}\,.$
@@
B.3 Proof of Theorem 5
Consider a single test point $\mathbf{x}_{*}$. Fix $m=1$. Recall that $H_{t}=\left(X_{\ell},Y_{\ell}\right)_{\ell\in[t-1]}$ is the history till round $t$. Then the posterior variance is
@@ B.3 Proof of Theorem 5
Now fix an $\mathbf{x}_{i}\in\mathcal{X}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Define $\hat{\sigma}_{t,i}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i}^{2}]=\sigma_{i}^{2}$. Then using Lemma 10 we can show that with probability $(1-\delta)$
$\sigma_{i}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\leq\hat{\sigma}_{t,i}^{2}\leq\sigma_{i}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]$ (5)
@@
$\frac{1}{2}\sigma^{2}_{i}\leq\frac{1}{m}$ $\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}\leq\frac{5}{2}\sigma^{2}_{i}.$
Hence, to minimize the quantity $\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2}$ we should be minimizing the variance $2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$. Observe that minimizing the variance in SAL leads to minimizing the quantity $2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$ which is same as minimizing the score $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}$ for GO. The claim of the theorem follows.
@@
Fix round $t\in[T]$, a sample point $\mathbf{x}_{i}$, and a test point $\mathbf{x}_{*}$, and failure probability $\delta\in(0,1)$. Let
$m=m_{t,i}-1$
and suppose that $m>4\log(1/\delta)$. Define $\hat{\sigma}_{t,i}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i}^{2}]=\sigma_{i}^{2}$. Then
$\mathbb{P}\left(\hat{\sigma}_{t,i}^{2}\leq\sigma_{i}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\right)\leq\delta$
@@
$\mathbb{P}\left(\hat{\sigma}_{t,i}^{2}\geq\sigma_{i}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\right)\leq\delta$

### v2

@@ B.2 Proof of Lemma 3
Now note that $\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}$ is a positive semi-definite matrix for any $\mathbf{x}_{j}$ and any positive semi-definite matrix $\mathbf{A}$. As a result, $\mathbf{x}_{*,k}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*,k}\geq 0$ holds for any $\mathbf{x}_{*,k}$ and thus
$f(S_{+})=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*,k}^{\top}-\mathbf{x}_{*,k}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*,k}\leq\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*,k}^{\top}=f(S)\,.$
@@
B.3 Proof of Theorem 1
The proof is under the assumption that at round $t$, the training examples can be partitioned as $\mathcal{X}_{\text{examples}}=S_{k}\cup\overline{S}_{k}$. The set $S_{k}$ represents examples that are close to $\mathbf{x}_{*,k}$. The set $S_{k}$ is convex and define $\alpha_{k}\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\geq\alpha_{k}$ for all $\mathbf{x},\mathbf{y}\in S_{k}$. The set $\overline{S}_{k}$ represents examples that are not close to $\mathbf{x}_{*,k}$. It is defined $\beta_{k}\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta_{k}$ for all $\mathbf{x}\in S_{k}$ and $\mathbf{y}\in\overline{S}_{k}$. Define $\alpha_{\min}=\min_{k}\alpha_{k}$, and $\beta_{\max}=\beta_{\max}$.
Define the set $S=\cap_{k=1}^{K}S_{k}$ as the set of all examples that are close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$ and $\overline{S}=\cup_{k=1}^{K}\overline{S}_{k}$ as the set of all examples that are not close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$. Assume $S\neq\{\emptyset\}$ and $|S|>T$.
The inverse of the covariance matrix after $t-1$ observations is
$\widehat{\mathbf{\Sigma}}_{t}^{-1}=\mathbf{\Lambda}_{t}=\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}=\sum_{i=1}^{d}\lambda_{t,i}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\,.$
The latter is the eigendecompositon of $\mathbf{\Lambda}_{t}$, where $\lambda_{t,i}$ is the $i$-th largest eigenvalue and $\mathbf{v}_{t,i}$ is the corresponding eigenvector. Note that $\mathbf{\Lambda}_{t}^{-1}=\sum_{i=1}^{d}\lambda_{t,i}^{-1}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}$. To simplify exposition, we assume that all examples have unit length. We analyze the eigenvalues of $\mathbf{\Lambda}_{t}$ first.
@@ B.3 Proof of Theorem 1
For all $i\in[d]$, $1\leq\lambda_{t,i}\leq t$. Moreover, let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\lambda_{t,1}\geq\alpha_{\min}^{2}(t-1)+1$.
@@
The first claim follows directly from the definition of $\mathbf{\Lambda}_{t}$ and that $\|X_{\ell}\|_{2}\leq 1$. The second claim is proved using the definition of the maximum eigenvalue,
$\lambda_{t,1}=\mathbf{v}_{t,1}^{\top}\mathbf{\Lambda}_{t}\mathbf{v}_{t,1}\geq\mathbf{x}_{*,k}^{\top}\left(\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)\mathbf{x}_{*,k}\geq\alpha_{\min}^{2}(t-1)+1\,.$
The last inequality follows from $\mathbf{x}_{*,k}^{\top}X_{\ell}\geq\alpha_{\min}$. This completes the proof. ∎
@@
Let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\mathbf{v}_{t,1}\in S$. Moreover, let $\beta_{\max}\geq 1-\alpha_{\min}^{2}$. Then $\mathbf{v}_{t,i}\in\overline{S}$ for all $i\geq 2$.
@@
Since all $X_{\ell}\in S$ and $S$ is a convex set, $\mathbf{v}_{t,1}\in S$. Now take any $\mathbf{x}\in S$ and $i\geq 2$, and note that
$(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}\leq\sum_{i=2}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1-(\mathbf{x}^{\top}\mathbf{v}_{t,1})^{2}\leq 1-\alpha_{\min}^{2}\,.$
We use that $\sum_{i=1}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1$ and $\mathbf{x}^{\top}\mathbf{v}_{t,1}\geq\alpha_{\min}$. Therefore, when $\beta_{\max}\geq 1-\alpha_{\min}^{2}$, we have $\mathbf{v}_{t,i}\in\overline{S}$ for all $i\geq 2$. ∎
Our analysis has two parts. First, we bound the approximation error under the assumption that $X_{\ell}\in S$ holds for all $\ell\in[T]$. Second, we show how to choose $\alpha_{\min}$ and $\beta_{\max}$ to guarantee this. We start with the approximation error.
@@
Let $X_{\ell}\in S$ hold for all $\ell\in[T]$. Then for any $\mathbf{x}_{*,k}$
$\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*,k}\leq\frac{1}{\alpha_{\min}^{2}T+1}+(1-\alpha_{\min}^{2})\,.$
@@
$\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*,k}=\sum_{i=1}^{d}\lambda_{T+1,i}^{-1}\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i}\mathbf{v}_{T+1,i}^{\top}\mathbf{x}_{*,k}\leq\frac{(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,1})^{2}}{\alpha^{2}_{\min}T+1}+\sum_{i=2}^{d}(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i})^{2}\,.$
The inequality uses lower bounds in Lemma 6. Then we apply $\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,1}\leq 1$ and $\sum_{i=2}^{d}(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i})^{2}\leq 1-\alpha^{2}_{\min}$. ∎
Now we prove by induction that $X_{\ell}\in S$ holds for all $\ell\in[T]$.
@@
Let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Suppose that $t\leq\frac{\alpha^{2}_{\min}}{(\beta+\sqrt{2})\beta d}$. Then $\mathbf{x}_{t}\in S$.
@@
Our algorithm chooses a example $\mathbf{x}_{t}\in S$ when for all $\mathbf{x}_{*,k}$
$\frac{\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}}{1+\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}}\geq\frac{\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}}{1+\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}}$
holds for any $\mathbf{x}\in S$ and $\mathbf{y}\in\overline{S}$. Since $0\leq\mathbf{v}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{v}\leq 1$ for any $\|\mathbf{v}\|_{2}\leq 1$, the above event occurs when
$\min_{k}(\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x})^{2}=\min_{k}\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}\geq 2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}=2\max_{k\in[K]}(\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y})^{2}\,.$
@@
$\max_{k\in[K]}|\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}|\leq\max_{k\in[K]}\sum_{i=1}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{y}|\leq\beta_{\max}d\,.$
Here we use $\lambda_{t,i}\geq 1$ (Lemma 6), and that $\mathbf{v}_{t,1}^{\top}\mathbf{y}\leq\beta_{\max}$ and $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta_{\max}$ when $i\geq 2$.
@@
$\min_{k}|\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}|\geq\min_{k}\lambda_{t,1}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,1}\mathbf{v}_{t,1}^{\top}\mathbf{x}|-\sum_{i=2}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{x}|$ $\geq\frac{\alpha^{2}_{\min}}{t}-\beta^{2}_{\max}d\,.$
To bound the first term, we use $\lambda_{t,1}\leq t$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\geq\alpha_{\min}$ and $\mathbf{v}_{t,1}^{\top}\mathbf{x}\geq\alpha_{\min}$. To bound the second term, we use $\lambda_{t,i}\geq 1$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta_{\max}$ and $\mathbf{v}_{t,i}^{\top}\mathbf{x}\leq\beta_{\max}$.
Now we chain all inequalities and get that our algorithm chooses a example $\mathbf{x}_{t}\in S$ when
$t\leq\frac{\alpha^{2}_{\min}}{(\beta_{\max}+\sqrt{2})\beta_{\max}d}\,.$
@@
B.4 Proof of Theorem 2
Consider the test examples $\{\mathbf{x}_{*}\}_{k=1}^{K}$. Recall that $H_{t}=\left(X_{\ell},Y_{\ell}\right)_{\ell\in[t-1]}$ is the history till round $t$. Then the posterior variance is
@@ B.4 Proof of Theorem 2
Now fix an $\mathbf{x}_{i}\in\mathcal{X}_{\text{examples}}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$ such that
$\widehat{\mathbf{\Sigma}}_{t,i}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)\right)^{-1}.$
Define $\hat{\sigma}_{t,i,k}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]=\sigma_{t,i,k}^{2}$. Then define $\sigma_{t,i,\max}^{2}=\max_{k\in[K]}\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]$. Then using Lemma 10 we can show that with probability $(1-\delta)$
$\sigma_{t,i,\max}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]$ $\leq\max_{k\in[K]}\hat{\sigma}_{t,i,k}^{2}\leq\sigma_{t,i,\max}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]$ (4)
@@
$\frac{1}{2}\sigma^{2}_{t,i,\max}\leq\max_{k\in[K]}\frac{1}{m}$ $\sum_{j=1}^{m}(\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k})^{2}\leq\frac{5}{2}\sigma^{2}_{t,i,\max}.$
Hence, to minimize the quantity $\max_{k\in[K]}\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}$ we should be minimizing the variance $2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}$. Observe that minimizing the variance in SAL leads to minimizing the quantity $2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}$ which is same as minimizing the score $\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}$ for GO. The claim of the theorem follows.
@@
Fix round $t\in[T]$, a sample example $\mathbf{x}_{i}$, and a test example $\mathbf{x}_{*,k}$, and failure probability $\delta\in(0,1)$. Suppose that $m>4\log(1/\delta)$. Define $\hat{\sigma}_{t,i,k}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]=\sigma_{i,k}^{2}$. Then
$\mathbb{P}\left(\hat{\sigma}_{t,i,k}^{2}\leq\sigma_{i,k}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\right)\leq\delta$
@@
$\mathbb{P}\left(\hat{\sigma}_{t,i,k}^{2}\geq\sigma_{i,k}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\right)\leq\delta$